OpenStudy (anonymous):

Solve the equation. Be sure to show your work. All final answers must be in simplest form. p+2/3=6

OpenStudy (angela210793):

let's do this together... subtract 2/3 both sides and u'll get p

OpenStudy (lgbasallote):

but what if it was actually \[\huge \frac{p+2}{3} = 6\]

OpenStudy (lgbasallote):

...then you're screwed :p

OpenStudy (angela210793):

Then...I was waiting for Sharron to correct me :P

OpenStudy (anonymous):

6-2/3=5 1/3

OpenStudy (anonymous):

p=5 1/3??????

OpenStudy (angela210793):

yea

OpenStudy (anonymous):

thank you